OpenStudy (anonymous):

I have another one: Find the exact value of log4 36- log4 9?

Parth (parthkohli):

\[\log (a) - \log(b) = \log\left(\frac{a}{b}\right)\]

OpenStudy (anonymous):

how would I plug it into the calculator when the base is 4? and the calculator has log base 10.

Parth (parthkohli):

Step 1: Throw the calculator into the dustbin.

Parth (parthkohli):

Step 2: Apply the identity I gave above. It works for all bases.

OpenStudy (dan815):

lol

OpenStudy (anonymous):

\[\log_{4} (36/9) \] \[\log_{4} 4 \] 1

OpenStudy (dan815):

step 3: beat up parth